Evaluate the following integral using Stokes theorem

Integralof over C (x+z)dx+(x-y)dy+xdz)

Where C is an ellipse

(x^2/16)+(y^2/25)=1 z=4
1
Expert's answer
2020-07-09T19:33:39-0400

We have "I=\oint\limits_C(x+z)dx+(x-y)dy+xdz".


The Stokes' formula is:


"\oint\limits_C(Pdx + Qdy + Rdz)=\iint\limits_S(\frac{\partial{R}}{\partial{y}}-\frac{\partial{Q}}{\partial{z}})dydz+"

"+(\frac{\partial{P}}{\partial{z}}-\frac{\partial{R}}{\partial{x}})dzdx + (\frac{\partial{Q}}{\partial{x}}-\frac{\partial{P}}{\partial{y}})dxdy".


Where "S -" the part of the plane "(z=4)" bounded by an ellipse "C".

And "C" is the ellipse "(\frac{x^2}{16}+\frac{y^2}{25} = \frac{x^2}{4^2}+\frac{y^2}{5^2}=\frac{x^2}{a^2}+\frac{y^2}{b^2})".

And "P = x+z", "Q=x-y", "R=x".


Find the necessary partial derivatives:

"\frac{\partial{R}}{\partial{y}}=0,
\frac{\partial{Q}}{\partial{z}}=0,
\frac{\partial{P}}{\partial{z}}=1,
\frac{\partial{R}}{\partial{x}}=1,
\frac{\partial{Q}}{\partial{x}}=1,
\frac{\partial{P}}{\partial{y}}=0."


So, we can write:

"\iint\limits_S(0-0)dydz+\iint\limits_S(1-1)dzdx+\iint\limits_S(1-0)dxdy".

"\iint\limits_Sdxdy = I".


The integrand is equal to one "(f(x, y)=1)", so the double integral is equal to the surface area "S".

It means that the "I" is equal to the area of the ellipse:

"I=\pi ab=4\times5\times\pi=20\pi."


Answer: "I=20\pi \approx62.8"